# A Smooth Approximation Framework for Weakly Convex Optimization

### v1

@@ 7 Numerical experiments
$\min_{\mathbf{x}\in\mathcal{X}\subseteq\mathbb{R}^{n}}f(\mathbf{x})\coloneqq\textstyle\frac{1}{m}\sum_{i=1}^{m}f(\mathbf{x},\xi_{i}),$
@@
  • Moreau envelop smoothing takes $f_{\eta}(\mathbf{x},\xi)=\operatornamewithlimits{arg\,min}_{\mathbf{y}}\{f(\mathbf{y},\xi)+\frac{2\rho+\eta^{-1}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\}$. Experiments using Moreau envelope use $h(z)=z^{2}$ to ensure a closed-form solution [davis2019stochastic].
@@ 7.1 Experiment setups
Dataset generation [gao2024stochastic]
Let $A\in\mathbb{R}^{m\times n}$ have $\{\mathbf{a}_{i}\}$ as its rows. Given a condition number parameter $\kappa\geq 1$, we generate $A=QD\in\mathbb{R}^{m\times n}$, where each element of $Q\in\mathbb{R}^{m\times n}$ is sampled from standard normal distribution and $D\in\mathbb{R}^{n\times n}$ is a diagonal matrix whose diagonal elements are evenly distributed between 1 and $\kappa$. We generate $\mathbf{x}^{\star}\sim\mathcal{N}(0,I_{n})$ and let $b_{i}=h(\langle\mathbf{a}_{i},\mathbf{x}^{\star}\rangle)+\theta_{i}\varepsilon_{i}$, where $\varepsilon_{i}$ simulates corruption by random noise, $\theta_{i}\sim\operatorname{Bernoulli}(p)$ and $\varepsilon_{i}\sim\mathcal{N}(0,25)$. Here $p\in[0,1]$ represents the fraction of corrupted data on expectation.
@@
    Dataset. We use $m=300,n=100$ to test deterministic algorithms and $m=1000,n=20$ to test stochastic algorithms. In particular, we use $m=30,n=10$ when Moreau envelope smoothing is used.
@@
    Initial point. We set the initial point of all the algorithms to be $\mathbf{x}^{0}\sim\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|}$, where $\hat{\mathbf{x}}\sim\mathcal{N}(0,I_{n})$.
@@ 7.2 Experiments on deterministic problems
Figure 1: Deterministic problems. First row: $h(z)=z^{2}$; Second row: $h(z)=z^{5}+z^{3}+1$; Within each row from left to right: $(\kappa,p)\in\{(1,0),(10,0),(1,0.2),(10,0.2)\}$. x-axis: iteration number. y-axis: $f(\mathbf{x}^{k})$ .
@@
Range of optimal stepsize for different smoothing approaches
Our theory suggests that Moreau envelope smoothing has a different range of stepsizes from Nesterov smoothing. Our second experiment investigates this behavior. Figure 2 illustrates that Moreau envelope smoothing typically admits larger stepsizes than Nesterov smoothing. This observation aligns with our theoretical findings.
@@ 7.3 Experiments on stochastic problems
Figure 3 illustrates the performance of different stochastic algorithms on the tested problems. We observe that although smoothing does not always yield faster convergence, it does lead to more robust convergence across different instances.
Robustness of Moreau envelope smoothing
Since Moreau envelope smoothing essentially uses the direction provided by stochastic proximal point, we expect it to inherit the stability properties of the proximal point method [deng2021minibatch, asi2019stochastic]. Figure 4 illustrates the number of iterations an algorithm takes before convergence for different values of $\alpha_{0}$. It is observed that Moreau envelope smooth is indeed more robust than SGM and Nesterov smoothing.
@@ 7.4 Experiments on generalized Lipschitz problems
This section conducts additional experiments to demonstrate the performance of AGLS (Algorithm 3). In particular, we consider the piecewise quadratic
$\min_{\mathbf{x}\in\mathbb{R}^{n}}\max_{1\leq j\leq m}\{\tfrac{1}{2}\langle\mathbf{x},A_{j}\mathbf{x}\rangle-\langle\mathbf{b}_{j},\mathbf{x}\rangle\},$
where $\{A_{i}\}$ are symmetric positive definite matrices.
@@
We generate $A_{i}=CC^{\top}$ with $C_{ij}\sim\mathcal{N}(0,1)$ and $\mathbf{b}\sim\mathcal{N}(0,I_{n})$. We generate $\{A_{i}\}$ such that at least one of them is positive definite, so that the objective function is coercive. We test $m\in\{5,10\}$ and $n\in\{20,100\}$.
@@
We compare AGLS applied to the softmax-smoothed version of the problem with normalized gradient descent (NGD) that is designed for generalized Lipschitz problems [grimmer2019convergence]. $\mathbf{x}^{k+1}=\mathbf{x}^{k}-\eta\frac{f^{\prime}(\mathbf{x}^{k})}{\|f^{\prime}(\mathbf{x}^{k})\|}$. As Figure 5 shows, our algorithm demonstrates competitive performance in practice.
Figure 5: Experiments on the comparison between accelerated gradient descent with line-search and normalized gradient descent for generalized smooth problems. From left to right: $(m,n)\in\{(5,20),(5,100),(10,20),(10,100)\}$. x-axis: iteration number. y-axis: $f(\mathbf{x}^{k})$ .
@@ Appendix A Auxiliary results
A.1 Local Lipschitzness and subgradient bound
@@
First, we show $b\Rightarrow a$”. Since $g$ is weakly convex on $\mathcal{S}$, it is locally Lipschitz continuous (see cui2021modern). By the mean value theorem [mordukhovich2022convex, Theorem 7.44], we have
@@
for some $\zeta\in\partial f(\theta\mathbf{x}+(1-\theta)\mathbf{y})$, $\theta\in(0,1)$. Applying Cauchy-Schwarz’s inequality, we immediately obtain part a.
Next, we show $a\Rightarrow b$”. The idea follows closely the proof of a similar result for a convex function [beck2017first, Theorem 3.61]. Since $\mathcal{S}$ is an open set, for any $\mathbf{x}\in\mathcal{S}$, $\mathbf{v}\in\partial g(\mathbf{x})$, there exists an $\bar{\varepsilon}\in(0,\infty)$ such that for any $\varepsilon\in(0,\bar{\varepsilon})$, the neighborhood $N_{\varepsilon}(\mathbf{x})=\{\mathbf{y}:\lVert\mathbf{y}-\mathbf{x}\rVert\leq\varepsilon\}\subseteq\mathcal{S}$. Moreover, let $\mathbf{v}^{\dagger}\in\mathbb{E}$ be a vector such that $\lVert\mathbf{v}^{\dagger}\rVert=1$, and $\langle\mathbf{v},\mathbf{v}^{\dagger}\rangle=\lVert\mathbf{v}\rVert_{*}$. Therefore, $\mathbf{x}+\varepsilon\mathbf{v}^{\dagger}\in\mathcal{S}$. Using Lipschitz continuity and weak convexity (of modulus $\mu$, $\mu>0$), we have
@@
Dividing both sides by $\varepsilon$ gives $\lVert\mathbf{v}\rVert-\frac{\mu}{2}\varepsilon\leq L.$ Since $\varepsilon$ can be arbitrarily small, we conclude that $\lVert\mathbf{v}\rVert\leq L$, thereby completing the proof.
A.2 Three-point inequality
@@
Let $f$ be a $\rho$-weakly convex function and define $\mathbf{y}_{\mathbf{x}}=\mathrm{prox}_{f/\beta}(\mathbf{x})$. Then, for any $\mathbf{z}\in\mathbb{R}^{d}$,
@@
If $f$ is $\mu$-strongly convex, then the inequality (61) holds with $\rho=-\mu$.
@@ Appendix B Missing Proofs
Summing up the above two inequalities, we obtain
Applying Cauchy-Schwartz inequality, we immediately obtain
@@
$=\beta\langle\mathbf{x}-\mathbf{y}_{\mathbf{x}},\hat{\mathbf{x}}-\mathbf{x}\rangle+\frac{\beta}{2}\lVert\mathbf{y}_{\hat{\mathbf{x}}}-\hat{\mathbf{x}}-(\mathbf{y}_{\mathbf{x}}-\mathbf{x})\rVert^{2}-\frac{\rho}{2}\lVert\mathbf{y}_{\mathbf{x}}-\mathbf{y}_{\hat{\mathbf{x}}}\rVert^{2}$
In view of (65) and (66), we have that
@@
$\geq-\frac{\beta-\rho}{2}\lVert\mathbf{y}_{\mathbf{x}}-\mathbf{y}_{\hat{\mathbf{x}}}\rVert^{2}+\frac{\beta}{2}\lVert\mathbf{x}-\hat{\mathbf{x}}\rVert^{2}\geq-\frac{\rho}{2(1-\rho/\beta)}\lVert\mathbf{x}-\hat{\mathbf{x}}\rVert^{2}$
@@
$|g^{\prime}(x)-g^{\prime}(y)|$ $=|x^{p-1}-y^{p-1}-\rho(x-y)|=\left|\sum_{k=0}^{p-2}x^{p-2-k}y^{k}-\rho\right||x-y|$ (71)
Thus, $g$ is $\mathcal{L}_{g}$-generalized smooth with $\mathcal{L}_{g}(x,y)=\left|\sum_{k=0}^{p-2}x^{p-2-k}y^{k}-\rho\right|$.
@@
Part 1). In view of (68) and (69), we conclude the quadratic bound (26). The Lipschitz smoothness follow from a more general result [boob2024level, Lemma 7]. Here, we use the argument of Corollary 3.4 [hoheisel2010proximal] to give a self-contained proof specified for the Moreau envelope. Note that
@@
$=\Big(\frac{\rho}{1-\rho/\beta}\Big)^{2}\lVert\mathbf{x}-\hat{\mathbf{x}}\rVert^{2}$
@@
For Part 2), placing $\mathbf{z}=\mathbf{x}$ in (61) gives $f(\mathbf{x})-f^{\beta}(\mathbf{x})\geq\frac{\beta-\rho}{2}\lVert\mathbf{x}-\mathbf{y}_{\mathbf{x}}\rVert^{2}=\frac{(1-\rho/\beta)}{2\beta}\lVert\nabla f^{\beta}(\mathbf{x})\rVert^{2}.$ Let $\mathbf{v}\in\partial f(\mathbf{x})$ denote a subgradient. Using the definition of $f^{\beta}(\mathbf{x})$, we have
@@
$=f(\mathbf{x})-\frac{\lVert\mathbf{v}\rVert^{2}}{\beta-\rho}\leq f(\mathbf{x})-\frac{\|\partial f(\mathbf{x})\|^{2}}{\beta-\rho},$
where the last inequality holds since $\lVert\partial f(\mathbf{x})\rVert\leq\lVert\mathbf{v}\rVert$.
@@
$\langle\nabla f_{\eta}(\mathbf{x}^{k-1}),\hat{\mathbf{x}}^{k}-\mathbf{x}\rangle+r(\hat{\mathbf{x}}^{k})-r(\mathbf{x})\leq\frac{1}{2\gamma}(\lVert\mathbf{x}-\mathbf{x}^{k-1}\rVert^{2}-\lVert\hat{\mathbf{x}}^{k}-\mathbf{x}^{k-1}\rVert^{2}-\lVert\hat{\mathbf{x}}^{k}-\mathbf{x}\rVert^{2}).$ (77)
@@
$\phi_{\eta}(\mathbf{x}^{k})$ $\leq\phi_{\eta}(\mathbf{x}^{k-1})+\frac{1}{2L_{\eta}}\lVert\bm{\delta}^{k-1}\rVert^{2}-\frac{1-L_{\eta}\gamma}{\gamma}\lVert\mathbf{x}^{k}-\mathbf{x}^{k-1}\rVert^{2}.$
@@
$\lVert G^{k}\rVert^{2}\leq\frac{1}{\gamma-L_{\eta}\gamma^{2}}[\phi_{\eta}(\mathbf{x}^{k-1})-\phi_{\eta}(\mathbf{x}^{k})]+\frac{1}{(2\gamma-2L_{\eta}\gamma^{2})L_{\eta}}\lVert\bm{\delta}^{k-1}\rVert^{2}.$ (79)
@@
Combining (79) and (80), we obtain
$\lVert\hat{G}^{k}\rVert^{2}\leq\frac{2}{\gamma-L_{\eta}\gamma^{2}}[\phi_{\eta}(\mathbf{x}^{k-1})-\phi_{\eta}(\mathbf{x}^{k})]+\frac{1}{(\gamma-L_{\eta}\gamma^{2})L_{\eta}}\lVert\bm{\delta}^{k-1}\rVert^{2}+2\lVert\bm{\delta}^{k-1}\rVert^{2}$
@@
Sum up over $k=1,2,\ldots,$$K$, take expectation over all the randomness, and notice that $\mathbb{E}[\lVert\bm{\delta}^{k}\rVert^{2}]\leq\frac{\sigma^{2}}{m}$:
$\frac{1}{K}\sum^{K}_{k=1}\mathbb{E}[\lVert\hat{G}^{k}\rVert^{2}]$ $\leq\frac{8L_{\eta}\mathbb{E}[\phi_{\eta}(\mathbf{x}^{0})-\phi_{\eta}(\mathbf{x}^{K})]}{K}+\frac{6\sigma^{2}}{m}$
@@
where the second inequality follows from $\phi_{\eta}(\mathbf{x})\leq\phi(\mathbf{x})$ in the definition of smooth approximation and the last inequality uses $\Delta\geq\phi(\mathbf{x}^{0})-\min_{\mathbf{x}}\phi(\mathbf{x})$.
@@
The key idea follows from the Moreau envelope-based analysis in davis2019stochastic. We set $\hat{\mathbf{x}}^{k}=\mathrm{prox}_{\phi_{\eta}/\hat{\rho}}(\mathbf{x}^{k-1})$. By the optimality condition of the proximal mapping, we have
@@
$\leq{}-\frac{\gamma^{-1}-\hat{\rho}}{2}\lVert\mathbf{x}^{k}-\mathbf{x}^{k-1}\rVert^{2}-\frac{\hat{\rho}-\bar{\rho}+\gamma^{-1}}{2}\lVert\mathbf{x}^{k}-\hat{\mathbf{x}}^{k}\rVert^{2}+\frac{\gamma^{-1}-\hat{\rho}}{2}\lVert\hat{\mathbf{x}}^{k}-\mathbf{x}^{k-1}\rVert^{2}.$ (82)
@@
$=f_{\eta}(\hat{\mathbf{x}}^{k})-f_{\eta}(\mathbf{x}^{k})+\langle\nabla f_{\eta}(\mathbf{x}^{k-1}),\mathbf{x}^{k}-\hat{\mathbf{x}}^{k}\rangle+\langle\bm{\delta}^{k-1},\mathbf{x}^{k}-\hat{\mathbf{x}}^{k}\rangle$
@@
$\geq-\frac{\bar{\rho}}{2}\lVert\mathbf{x}^{k-1}-\hat{\mathbf{x}}^{k}\rVert^{2}-\frac{L_{\eta}}{2}\lVert\mathbf{x}^{k}-\mathbf{x}^{k-1}\rVert^{2}+\langle\bm{\delta}^{k-1},\mathbf{x}^{k}-\hat{\mathbf{x}}^{k}\rangle,$ (83)
where (82) is due to the $\bar{\rho}$-weak convexity and $L_{\eta}$-smoothness of $f_{\eta}(\mathbf{x})$. Combining (82) and (83),
@@
$\quad-\frac{\hat{\rho}-\rho+\gamma^{-1}}{2}\lVert\mathbf{x}^{k}-\hat{\mathbf{x}}^{k}\rVert^{2}+\frac{\gamma^{-1}+\bar{\rho}-\hat{\rho}}{2}\lVert\hat{\mathbf{x}}^{k}-\mathbf{x}^{k-1}\rVert^{2}.$

### v3

@@ 7 Numerical experiments
$\min_{\mathbf{x}\in\mathcal{X}\subseteq\mathbb{R}^{d}}f(\mathbf{x})\coloneqq\textstyle\frac{1}{m}\sum_{i=1}^{m}f(\mathbf{x},\xi_{i}),$
@@
  • Moreau envelop smoothing takes $f_{\eta}(\mathbf{x},\xi)=\operatornamewithlimits{arg\,min}_{\mathbf{y}}\{f(\mathbf{y},\xi)+\frac{2\rho+\eta^{-1}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\}$. Experiments using Moreau envelope use $h(z)=z^{2}$ to ensure a closed-form solution [13].
@@ 7.1 Experiment setups
Dataset generation [18]
Let $A\in\mathbb{R}^{m\times d}$ have $\{\mathbf{a}_{i}\}$ as its rows. Given a condition number parameter $\kappa\geq 1$, we generate $A=QD\in\mathbb{R}^{m\times d}$, where each element of $Q\in\mathbb{R}^{m\times d}$ is sampled from standard normal distribution and $D\in\mathbb{R}^{d\times d}$ is a diagonal matrix whose diagonal elements are evenly distributed between 1 and $\kappa$. We generate $\mathbf{x}^{\star}\sim\mathcal{N}(0,I_{d})$ and let $b_{i}=h(\langle\mathbf{a}_{i},\mathbf{x}^{\star}\rangle)+\theta_{i}\varepsilon_{i}$, where $\varepsilon_{i}$ simulates corruption by random noise, $\theta_{i}\sim\operatorname{Bernoulli}(p)$ and $\varepsilon_{i}\sim\mathcal{N}(0,25)$. Here $p\in[0,1]$ represents the fraction of corrupted data on expectation.
@@
    Dataset. We use $m=300,d=100$ to test deterministic algorithms and $m=1000,d=20$ to test stochastic algorithms. In particular, we use $m=30,d=10$ when Moreau envelope smoothing is used.
@@
    Initial point. We set the initial point of all the algorithms to be $\mathbf{x}^{0}\sim\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|}$, where $\hat{\mathbf{x}}\sim\mathcal{N}(0,I_{d})$.
@@ 7.2 Experiments on deterministic problems
Figure 1: Deterministic problems. First row: $h(z)=z^{2}$; Second row: $h(z)=e^{x}+10$; Within each row from left to right: $(\kappa,p)\in\{(1,0),(10,0),(1,0.2),(10,0.2)\}$. x-axis: iteration number. y-axis: $f(\mathbf{x}^{k})$ .
@@
Range of Optimal Stepsizes for Different Smoothing Approaches.
Our theory predicts that Moreau envelope smoothing admits a different range of admissible stepsizes than Nesterov smoothing. To examine this behavior empirically, our second experiment studies the sensitivity of each method to the choice of stepsize. Figure 2 reports the number of iterations required to achieve convergence under various stepsize configurations. As the figure illustrates, the admissible stepsizes for Moreau envelope smoothing are larger than for Nesterov smoothing, which is consistent with our theoretical findings.
@@ 7.3 Experiments on stochastic problems
Figure 3 summarizes the performance of the stochastic algorithms on the tested problems. With appropriate choices of the smoothing parameter and the minibatch size for the proximal-point subproblems, we observe that smoothing often achieves significantly better performance than the vanilla SGM. Although smoothing does not universally accelerate convergence, it typically yields more robust behavior across different instances. Notably, on certain problems, convergence is observed only when AGLS-SIPP is employed.
Robustness of Moreau Envelope Smoothing.
Moreau envelope smoothing exploits the full stochastic objective via a proximal update, rather than relying on a first-order approximation as in gradient-based methods. As a result, it is closely related to the stochastic proximal point method, and we therefore expect it to inherit the stability properties of proximal algorithms [15, 1]. Figure 4 reports the number of iterations required to meet the stopping criterion for different values of $\alpha_{0}$. The results confirm that Moreau envelope smoothing is indeed more robust than both SGM and Nesterov smoothing, consistent with the existing theoretical findings.
@@ 7.4 Experiments on generalized Lipschitz problems
This section reports additional experiments evaluating our smoothing algorithm on non-Lipschitz problems. In particular, we consider the piecewise quadratic problem
$\min_{\mathbf{x}\in\mathbb{R}^{d}}\max_{1\leq j\leq m}\{\tfrac{1}{2}\langle\mathbf{x},A_{j}\mathbf{x}\rangle-\langle\mathbf{b}_{j},\mathbf{x}\rangle\},$
where $\{A_{i}\}$ are symmetric matrices.
@@
We generate $A_{i}=CC^{\top}-\sigma I$ with $C_{ij}\sim\mathcal{N}(0,1),\mathbf{b}\sim\mathcal{N}(0,I_{d})$ and $\sigma\in\{0,1\}$. When $\sigma=0$, we generate $\{A_{i}\}$ such that at least one of them is positive definite, so that the objective function is coercive. We test $m\in\{5,10\}$ and $d\in\{20,100\}$. As a result, the objective is not globally Lipschitz, but it satisfies Assumption 4.
@@
When $\sigma=0$, we compare AGLS applied to the softmax-smoothed objective with the normalized (sub)gradient method (NSGM) that is designed for generalized Lipschitz problems [23]. The NSGM update takes the form $\mathbf{x}^{k+1}=\mathbf{x}^{k}-\eta\frac{f^{\prime}(\mathbf{x}^{k})}{\|f^{\prime}(\mathbf{x}^{k})\|}$. When $\sigma=1$, we use AGLS-SIPP as the benchmark algorithm. As Figure 5 illustrates, on non-Lipschitz optimization tasks, our method exhibits competitive performance in practice compared to the subgradient-based method.
Figure 5: Experiments on the comparison between AGLS (AGLS-SIPP) and NSGM for generalized smooth problems. The first row corresponds to $\sigma=0$ and the second row corresponds to $\sigma=1$. From left to right: $(m,d)\in\{(5,20),(5,100),(10,20),(10,100)\}$. $x$-axis: iteration number. $y$-axis: $f(\mathbf{x}^{k})$ (or $f(\mathbf{x}^{k})-f^{\star}$ for convex problems) .
@@ Appendix A Auxiliary results
Local Lipschitzness and subgradient bound
@@
First, we show $b\Rightarrow a$”. Since $g$ is weakly convex on $\mathcal{S}$, it is locally Lipschitz continuous (see Cui and Pang [12, Lemma 4.4.1]). By the mean value theorem [32, Theorem 7.44], we have
@@
for some $\zeta\in\partial g(\theta\mathbf{x}+(1-\theta)\mathbf{y})$, $\theta\in(0,1)$. Applying Cauchy-Schwarz’s inequality, we immediately obtain part a.
Next, we show $a\Rightarrow b$”. The idea follows closely the proof of a similar result for a convex function [3, Theorem 3.61]. Since $\mathcal{S}$ is an open set, for any $\mathbf{x}\in\mathcal{S}$, $\mathbf{v}\in\partial g(\mathbf{x})$, there exists an $\bar{\varepsilon}\in(0,\infty)$ such that for any $\varepsilon\in(0,\bar{\varepsilon})$, the neighborhood $N_{\varepsilon}(\mathbf{x})=\{\mathbf{y}:\lVert\mathbf{y}-\mathbf{x}\rVert\leq\varepsilon\}\subseteq\mathcal{S}$. Moreover, let $\mathbf{v}^{\dagger}\in\mathbb{E}$ be a vector such that $\lVert\mathbf{v}^{\dagger}\rVert=1$, and $\langle\mathbf{v},\mathbf{v}^{\dagger}\rangle=\lVert\mathbf{v}\rVert_{*}$. Therefore, $\mathbf{x}+\varepsilon\mathbf{v}^{\dagger}\in\mathcal{S}$. Using Lipschitz continuity and weak convexity (of modulus $\mu$, $\mu>0$), we have
@@
Dividing both sides by $\varepsilon$ gives $\lVert\mathbf{v}\rVert-\frac{\mu}{2}\varepsilon\leq L.$ Since $\varepsilon$ can be arbitrarily small, we conclude that $\lVert\mathbf{v}\rVert\leq L$, thereby completing the proof. ∎
Three-point inequality
The following three-point property is known, and it will be used several times in our analysis.
@@
Let $f$ be a $\rho$-weakly convex function and define $\mathbf{y}_{\mathbf{x}}=\mathrm{prox}_{f/\beta}(\mathbf{x})$ for some $\beta\geq\rho$. Then, for any $\mathbf{z}\in\mathbb{R}^{d}$,
@@
If $f$ is $\mu$-strongly convex ($\mu>0$), then the inequality (60) holds with $\rho=-\mu$.
@@ Appendix B Missing Proofs
$=\beta\langle\mathbf{x}-\mathbf{y}_{\mathbf{x}},\hat{\mathbf{x}}-\mathbf{x}\rangle+\frac{\beta}{2}\lVert\mathbf{y}_{\hat{\mathbf{x}}}-\hat{\mathbf{x}}-(\mathbf{y}_{\mathbf{x}}-\mathbf{x})\rVert^{2}-\frac{\rho}{2}\lVert\mathbf{y}_{\mathbf{x}}-\mathbf{y}_{\hat{\mathbf{x}}}\rVert^{2}.$
It follows that
@@
$=\frac{\beta-\rho}{2}\big\|\mathbf{y}_{\mathbf{x}}-\mathbf{y}_{\hat{\mathbf{x}}}-\frac{\beta}{\beta-\rho}(\mathbf{x}-\hat{\mathbf{x}})\big\|^{2}-\frac{\rho}{2(1-\rho/\beta)}\lVert\mathbf{x}-\hat{\mathbf{x}}\rVert^{2}$
$\geq-\frac{\rho}{2(1-\rho/\beta)}\lVert\mathbf{x}-\hat{\mathbf{x}}\rVert^{2}.$
@@
$|g^{\prime}(x)-g^{\prime}(y)|$ $=|x^{p-1}-y^{p-1}-\rho(x-y)|=\big|\sum_{k=0}^{p-2}x^{p-2-k}y^{k}-\rho\big||x-y|$ (68)
Thus, $g$ is $\mathcal{L}_{g}$-generalized smooth with $\mathcal{L}_{g}(x,y)=\big|\sum_{k=0}^{p-2}x^{p-2-k}y^{k}-\rho\big|$.
@@
Part 1). In view of (65) and (66), we conclude the quadratic bound (25). The Lipschitz smoothness follow from a more general result [5, Lemma 7]. Here, we use the argument of Corollary 3.4 [24] to give a self-contained proof specified for the Moreau envelope. We first establish some nonexpansiveness properties. Summing up (62) and (63), we obtain
Applying Cauchy-Schwarz inequality, we immediately obtain
Note that
@@
$=\big(\frac{\rho}{1-\rho/\beta}\big)^{2}\lVert\mathbf{x}-\hat{\mathbf{x}}\rVert^{2},$
@@
For Part 2), placing $\mathbf{z}=\mathbf{x}$ in (60) gives $f(\mathbf{x})-f^{\beta}(\mathbf{x})\geq\frac{\beta-\rho}{2}\lVert\mathbf{x}-\mathbf{y}_{\mathbf{x}}\rVert^{2}=\frac{(1-\rho/\beta)}{2\beta}\lVert\nabla f^{\beta}(\mathbf{x})\rVert^{2}.$ Let $\mathbf{v}\in\partial f(\mathbf{x})$ denote a minimum-norm subgradient. Using the definition of $f^{\beta}(\mathbf{x})$, we have
@@
$=f(\mathbf{x})-\frac{\lVert\mathbf{v}\rVert^{2}}{2(\beta-\rho)}=f(\mathbf{x})-\frac{\|\partial f(\mathbf{x})\|^{2}}{2(\beta-\rho)}.$
@@
$\langle\nabla f_{\eta}(\mathbf{x}^{k-1}),\hat{\mathbf{x}}^{k}-\mathbf{x}\rangle+r(\hat{\mathbf{x}}^{k})-r(\mathbf{x})\leq\frac{1}{2\gamma}(\lVert\mathbf{x}-\mathbf{x}^{k-1}\rVert^{2}-\lVert\hat{\mathbf{x}}^{k}-\mathbf{x}^{k-1}\rVert^{2}-\lVert\hat{\mathbf{x}}^{k}-\mathbf{x}\rVert^{2}),$ (76)
@@
$\phi_{\eta}(\mathbf{x}^{k})$ $\leq\phi_{\eta}(\mathbf{x}^{k-1})+\frac{1}{2L_{\eta}}\lVert\bm{\delta}^{k-1}\rVert^{2}-\frac{1-L_{\eta}\gamma}{\gamma}\lVert\mathbf{x}^{k}-\mathbf{x}^{k-1}\rVert^{2}$
$\leq\phi_{\eta}(\mathbf{x}^{k-1})+\frac{1}{2L_{\eta}}\lVert\bm{\delta}^{k-1}\rVert^{2}-\frac{1-L_{\eta}\gamma}{\gamma}\lVert\mathbf{x}^{k}-\mathbf{x}^{k-1}\rVert^{2}.$
@@
$\lVert G^{k}\rVert^{2}\leq\frac{1}{\gamma-L_{\eta}\gamma^{2}}[\phi_{\eta}(\mathbf{x}^{k-1})-\phi_{\eta}(\mathbf{x}^{k})]+\frac{1}{(2\gamma-2L_{\eta}\gamma^{2})L_{\eta}}\lVert\bm{\delta}^{k-1}\rVert^{2}.$
@@
Combining (B.5) and (78), we obtain
$\lVert\hat{G}^{k}\rVert^{2}\leq\frac{2}{\gamma-L_{\eta}\gamma^{2}}[\phi_{\eta}(\mathbf{x}^{k-1})-\phi_{\eta}(\mathbf{x}^{k})]+\frac{1}{(\gamma-L_{\eta}\gamma^{2})L_{\eta}}\lVert\bm{\delta}^{k-1}\rVert^{2}+2\lVert\bm{\delta}^{k-1}\rVert^{2}.$
@@
Summing over $k=1,2,\ldots,K$, taking expectation with respect to all sources of randomness, and noting that $\mathbb{E}[\lVert\bm{\delta}^{k}\rVert^{2}]\leq\frac{\sigma^{2}}{m}$, we obtain
$\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}[\lVert\hat{G}^{k}\rVert^{2}]$ $\leq\frac{8L_{\eta}\mathbb{E}[\phi_{\eta}(\mathbf{x}^{0})-\phi_{\eta}(\mathbf{x}^{K})]}{K}+\frac{6\sigma^{2}}{m}$
@@
where the second inequality uses the fact that $\phi_{\eta}(\mathbf{x})\leq\phi(\mathbf{x})$, as established in the definition of the smooth approximation, and the final inequality follows from the bound $\Delta\geq\phi(\mathbf{x}^{0})-\min_{\mathbf{x}}\phi(\mathbf{x})$.
@@
The key idea follows from the Moreau envelope-based analysis in Davis and Drusvyatskiy [13]. We set $\hat{\mathbf{x}}^{k}=\mathrm{prox}_{\phi_{\eta}/\hat{\rho}}(\mathbf{x}^{k-1})$. By the optimality condition of the proximal mapping, we have
@@
$\leq{}-\frac{\gamma^{-1}-\hat{\rho}}{2}\lVert\mathbf{x}^{k}-\mathbf{x}^{k-1}\rVert^{2}-\frac{\hat{\rho}-\bar{\rho}+\gamma^{-1}}{2}\lVert\mathbf{x}^{k}-\hat{\mathbf{x}}^{k}\rVert^{2}+\frac{\gamma^{-1}-\hat{\rho}}{2}\lVert\hat{\mathbf{x}}^{k}-\mathbf{x}^{k-1}\rVert^{2}.$
@@
$\geq-\frac{\bar{\rho}}{2}\lVert\mathbf{x}^{k-1}-\hat{\mathbf{x}}^{k}\rVert^{2}-\frac{L_{\eta}}{2}\lVert\mathbf{x}^{k}-\mathbf{x}^{k-1}\rVert^{2}+\langle\bm{\delta}^{k-1},\mathbf{x}^{k}-\hat{\mathbf{x}}^{k}\rangle,$
where the last inequality is due to the $\bar{\rho}$-weak convexity and $L_{\eta}$-smoothness of $f_{\eta}(\mathbf{x})$. Combining the above two inequalities, we have
@@
$\quad-\frac{\hat{\rho}-\bar{\rho}+\gamma^{-1}}{2}\lVert\mathbf{x}^{k}-\hat{\mathbf{x}}^{k}\rVert^{2}+\frac{\gamma^{-1}+\bar{\rho}-\hat{\rho}}{2}\lVert\hat{\mathbf{x}}^{k}-\mathbf{x}^{k-1}\rVert^{2}.$